### v4

# CMAT: A Multi-Agent Collaboration Tuning Framework for Enhancing Small Language Models

Abstract
Open large language models (LLMs) have significantly advanced the field of natural language processing, showcasing impressive performance across various tasks. Despite the significant advancements in LLMs, their effective operation still relies heavily on human input to accurately guide the dialogue flow, with agent tuning being a crucial optimization technique that involves human adjustments to the model for better response to such guidance. Addressing this dependency, our work introduces the TinyAgent model, trained on a meticulously curated high-quality dataset. We also present the Collaborative Multi-Agent Tuning (CMAT) framework, an innovative system designed to augment language agent capabilities through adaptive weight updates based on environmental feedback. This framework fosters collaborative learning and real-time adaptation among multiple intelligent agents, enhancing their context-awareness and long-term memory. In this research, we propose a new communication agent framework that integrates multi-agent systems with environmental feedback mechanisms, offering a scalable method to explore cooperative behaviors. Notably, our TinyAgent-7B model exhibits performance on par with GPT-3.5, despite having fewer parameters, signifying a substantial improvement in the efficiency and effectiveness of LLMs 111Access models, demos at https://github.com/heimy2000/CMAT/tree/master.
1 Introduction
In the rapid development of the field of artificial intelligence, large language models (LLMs) such as BERT and GPT-4 OpenAI (2023) have become important cornerstones of natural language processing (NLP). These models utilize the Transformer architecture and effectively capture long-distance dependencies through multi-head self-attention mechanisms, demonstrating strong capabilities across various NLP tasks. With technological advancements, the performance and application scope of LLMs continue to expand, promising significant improvements in computational efficiency and functionality, including anticipated advanced features such as self-improvement, self-checking, and sparse expert models Liu et al. (2023).
However, it is noteworthy that the success of these models largely depends on human input to guide the correct dialogue. This dependency requires users to provide relevant and precise prompts based on their intentions and the feedback from the chat agent, raising a critical question: Can we replace human intervention with autonomous communication agents capable of steering conversations towards task completion with minimal human supervision?
Our research is driven by the need to overcome the significant challenges faced by LLMs in real-world deployments, particularly the high computational resource requirements, data biases, and lack of robustness. These issues limit their applicability in resource-constrained environments and highlight the urgency of enhancing model efficiency and adaptability Abid et al. (2021); Du et al. (2022). As demonstrated by Figure 1, we aim to address these limitations by optimizing models and training methods to enable smaller models to match the performance levels of larger models. Additionally, recognizing the potential of MAS to improve processing efficiency and system adaptability through agent cooperation, we seek to develop a collaborative agent framework. This framework aims to facilitate effective cooperation among agents, thereby overcoming the performance gap and propelling further research and innovation in the field of LLMs Ferry et al. (2018); Talwar et al. (2005). In our experiments, we evaluated the capabilities of large models with and without the use of prompts and observed that low-quality prompts can significantly degrade model performance. Consequently, we propose the Collaborative Multi-Agent Tuning (CMAT) framework.
The CMAT framework introduces a structured environment where individual agents, each with specialized roles and capabilities, work together to process information, make decisions, and solve complex tasks Hernández-Orallo et al. (2017). By sharing insights and learning from interactions within this multi-agent ecosystem, the framework allows for a more scalable and flexible approach to training LLMs Lewis et al. (2017). This collaborative effort not only helps in bridging the gap in performance between smaller and larger models but also fosters a more resilient system capable of adapting to new challenges without extensive human intervention Kaplan et al. (2020). Through CMAT, we aim to push the boundaries of what is possible with LLMs, making them more accessible and effective for a wider range of applications Rajpurkar et al. (2018).
The main contributions of our work are as follows:
We propose the CMAT framework which represents an innovative approach that allows for dynamic and real-time memory updates within multi-agent systems.
We design a novel role-playing mechanism for precise task allocation and enhanced agent communication, significantly boosting overall performance and cooperation.
We evaluated the fine-tuned TinyAgent models across multiple agent tasks, finding that in certain scenarios, their performance rivals that of advanced LLMs like GPT-4 and agentlm Zeng et al. (2023), demonstrating the potential efficiency and capabilities of compact models.
2 Related Work
2.1 LLMs Applications in a Multi-Agent Framework
We explore the applications of LLMs within multi-agent systems, highlighting their role versatility as users, assistants, and checkers, and their capability to offer bespoke support and solutions across such environments de Zarzà et al. (2023); Talebirad and Nadiri (2023). LLMs showcase remarkable adaptability to tasks through methods like supervised fine-tuning and real-time feedback learning, notably in tasks that require a sophisticated understanding and execution related to operating systems or databases Christianos et al. (2023); Li et al. (2023). Furthermore, LLMs are adept at enhancing communication and collaboration among agents, a critical component for addressing complex issues that necessitate multi-role coordination Zhao et al. (2021). Nevertheless, LLMs encounter specific challenges within multi-agent frameworks, especially in situations that demand a nuanced contextual comprehension and sustained memory retention, as well as adapting to fast-evolving environments and unforeseeable tasks Diallo et al. (2020). Issues such as data bias, security concerns, and the intricacies of crafting effective protocols for multi-agent cooperation stand as significant hurdles in this domain Zhang et al. (2017); García et al. (2015). Thus, by summarizing LLMs’ roles in multi-agent frameworks, we underscore the critical need for continued innovation and research exploration, aimed at overcoming these technological hurdles and leveraging the full potential of LLMs in complex systems Lu and Zhang (2020).
To enhance the adaptability and collaborative capabilities of LLMs in multi-agent systems, we’ve implemented memory modes, including long-term support and short-term memory with environmental feedback Liang et al. (2016). This allows LLMs to better interact, learn, and adapt in dynamic environments, leveraging past experiences and responding to changes swiftly.
2.2 The tuning method for LLMs
The main tuning methods include supervised fine-tuning and reinforcement learning Ouyang et al. (2022). Supervised fine-tuning enhances performance by training models on specific task datasets, and is especially suitable for tasks such as natural language understanding (NLU) Howard and Ruder (2018). On the other hand, reinforcement learning, guided by reward mechanisms, is suitable for handling complex and variable tasks Mnih et al. (2015). The effective combination of these two methods can significantly improve the performance of LLMs in various tasks. Notably, LLMs of reduced scale, such as those encompassing 1.8 billion parameters, can achieve performance levels akin to those of models with greater parameter counts, like 6 billion parameters, when supported by high-quality datasets Stiennon et al. (2020). This demonstrates that excellent data quality and appropriate tuning strategies play a decisive role in the performance of LLMs. Therefore, investing efforts in improving data quality and choosing the right tuning methods is essential for achieving optimal performance of LLMs in various application scenarios Howard and Ruder (2018). Through our work combining supervised fine-tuning with reinforcement learning, we’ve notably advanced LLM performance across a spectrum of tasks, showcasing significant improvements in task-specific benchmarks Ouyang et al. (2022).
3 Methodology
Our research is dedicated to the development and application of the Collaborative Multi-Agent Language Model Tuning (CMAT) framework. This framework aims to enhance the quality of decision-making, controllability, and operational efficiency within complex systems through the collaborative efforts of various agent roles, including the User, Assistant, and Checker. The CMAT framework employs advanced supervised fine-tuning techniques such as Low-Rank Adaptation (LoRA)Hu et al. (2021) and Prompt Tuning (PTuning)Lester et al. (2021), leveraging pre-trained datasets like AgentBench. Additionally, the framework incorporates interactive environment engagement and memory updates to further refine the model’s performance, drawing inspiration from the principles of Reinforcement Learning from Human Feedback (RLHF)Vázquez-Canteli and Nagy (2019).
3.1 Actor-Evaluator Dynamics in the CMAT Framework
In multi-agent systems, the Actor plays a crucial role in behavior generation through LLMs. The Actor responds to environmental state observations to generate text and actions, while also using reflection to assess and adjust its behavior Wang and Shi (2019). In traditional reinforcement learning (RL) frameworks, the Actor decides on actions based on a specific policy and receives information about the current state of the environment. In the CMAT framework, the Actor model utilizes approaches like Chain of Thought and ReAct, enhancing the traditional text generation models to delve into different facets of behavior generation.
3.2 Real-Time Review Loop and Interdisciplinary Research
Long-term Memory
Within the context of multi-agent systems, the significance of long-term memory for LLMs cannot be overstatedKato et al. (2022). It serves not merely as a storage for information but as an integrated knowledge management system, enabling LLMs to retain and access vital data over extended periodsBhadra (2022). Our research delves into interdisciplinary insights, particularly from psycholinguistics, to enhance the problem-solving capabilities of LLMs through the strategic utilization of Chain-of-Thought (CoT).
A pivotal discovery from our interdisciplinary exploration is the impact of CoT generation order on the correctness of LLMs’ responses. We have observed that when LLMs first generate a CoT before producing an answer, the correctness of the final answer is significantly higher compared to when the answer is generated first and then a CoT is constructed. Notably, the CoTs generated in both scenarios are identical, indicating that the order of generation is crucial. This suggests that while the models may have the potential to arrive at correct conclusions, their capabilities are not yet sufficient to do so autonomously. By employing a memory-assisted approach that prioritizes CoT generation, we can effectively mitigate this limitation and enhance the overall performance of LLMs in complex decision-making tasks.
Short-term Memory and Environmental Feedback
Conversely, short-term memory focuses on immediate contexts and the rapid processing of new information, which is essential for LLMs to adapt swiftly to dynamic environmentsPae et al. (2012); Liu and Guo (2019). This facet of memory is equally informed by psycholinguistic research, enabling LLMs to respond promptly to environmental changes and demands. Furthermore, environmental communication feedback mechanisms complement short-term memory by providing real-time inputs and interactions within the multi-agent frameworkYogatama et al. (2021). This continuous flow of information allows LLMs to adjust their strategies based on the latest data, facilitating a more agile and effective adaptation process.
By integrating the strengths of both long-term and short-term memory, informed by interdisciplinary research, our LLMs are better positioned to simulate human cognitive processes, offering a more human-like interaction experience. This cross-disciplinary approach not only deepens our understanding of memory system design but also paves the way for future AI research, optimizing LLMs’ performance and adaptability through the emulation of human thought processes.
| Model | BLEU-4 | ROUGE-1 | ROUGE-2 | ROUGE-L |
| codellama-7b | 25.01 | 45.91 | 29.83 | 26.24 |
| codellama-13b | 26.96 | 45.31 | 29.54 | 25.91 |
| tinyllama-1.8b | 43.38 | 59.86 | 37.81 | 42.86 |
The Reflexion Process
Artificial intelligence systems are increasingly adopting self-reflection mechanisms to enhance their decision-making capabilities. By systematically reviewing past actions and their outcomes, these systems can identify patterns and make informed adjustments to their strategies.
The CMAT framework introduces a Reflexion Process of collaboration between computation and cognition by initializing LLMs, users, assistants, checkers, and a set of tasks Li et al. (2018). Each task is executed by an assistant and verified by a checker. The assistant selects an action through interaction with LLMs, obtaining a new state and reward Che et al. (2021). If the checker verifies as correct, the system updates the state to and stores in memory mem; otherwise, it adjusts strategy and retries. The entire process not only involves confirmation of correct actions but also includes updating the strategy based on feedback from the checker, creating a continuous cycle of learning and adjustment. Each iteration aims to optimize the assistant’s decision-making strategy through practice and reflection, enhancing the efficiency and accuracy of task execution. Once all tasks are verified, it outputs the final states and evaluation results of all tasks Silver et al. (2017).Within this cycle, there is also a critical step: updating the checker’s strategy. This means that after each update of the strategy , not only does the assistant’s strategy need to be adjusted based on feedback, but the checker’s verification strategy also needs to be updated synchronously to ensure its decision-making logic is consistent with the latest environmental state and strategy. Such updates ensure the checker can accurately assess the assistant’s actions, further enhancing the system’s overall performance and reliability. Through this approach, the CMAT framework achieves a deeper level of collaboration and self-optimization between computation and cognition, enabling AI systems to face new challenges and environments with higher efficiency and adaptability.
Checker-In-The-Loop
To enhance the controllability of the role-playing framework, we have introduced an inspector agent that can inspect the role-playing Tao et al. (2023) agents for any issues. This enables a decision-making process for task solving similar to tree search. In practice, the inspector can be either an AI agent or a human.
3.3 Compared to Existing Practice
Our method stands out by dynamically updating memory through real-time interactions, enabling rapid adaptation to new experiences. It uses direct feedback from the environment for quick learning and incorporates both short-term and long-term memory updates for efficient decision-making. Our approach, enhanced with self-reflection and experience replay, offers deeper understanding and better response to complex scenarios, leading to more precise and rational decisions in changing conditions.
| LLM Type | Models | VER | OS | DB | KG | ALF | WS | M2W |
| API | gpt-3.5-turbo | 0613 | 31.6 | 15.7 | 25.9 | 16.0 | 64.1 | 16.0 |
| gpt-4 | 0613 | 42.4 | 32.0 | 58.8 | 78.0 | 61.6 | 29.0 | |
| text-davinci-003 | - | 20.1 | 16.3 | 34.9 | 20.0 | 61.7 | 26.0 | |
| text-davinci-002 | - | 8.3 | 16.7 | 41.5 | 16.0 | 56.3 | 9.0 | |
| OSS | tinyllama-1.1b Zhang et al. (2024) | - | 2.8 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| opt-1.3b Zhang et al. (2022) | - | 0.7 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | |
| opt-2.7b | - | 1.4 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | |
| qwen-1.8b | chat | 10.4 | 22.67 | 6.8 | 0.0 | 26.6 | 5.0 | |
| chatglm2-6b 222https://github.com/thudm/chatglm2-6b | v1.1 | 4.2 | 1.3 | 0.0 | 0.0 | 0.0 | 0.0 | |
| codellama-7b | instruct | 9.7 | 2.7 | 0.0 | 0.0 | 14.3 | 5.0 | |
| llama2-7b Touvron et al. (2023) | chat | 0.0 | 4.2 | 8.0 | 0.0 | 11.6 | 7.0 | |
| zephyr-7b Tunstall et al. (2023) | alpha | 12.5 | 9.7 | 5.0 | 8.0 | 45.0 | 11.0 | |
| baichuan2-6b Yang et al. (2023) | chat | 2.8 | 9.7 | 0.0 | 0.0 | 6.1 | 11.0 | |
| mpt-7b 333https://github.com/mosaicml/llm-foundry/ | chat | 5.6 | 9.7 | 12.7 | 0.0 | 0.0 | 0.0 | |
| qwen-7b | chat | 12.5 | 13.0 | 7.0 | 34.3 | 0.0 | 0.0 | |
| agentlm-7b | chat | 14.6 | 33.0 | 9.0 | 16.4 | 18.4 | 10.0 | |
| agentlm-7b(SFT) | chat | 17.4 | 37.0 | 10.0 | 17.4 | 26.6 | 10.0 | |
| tinyagent-1.8b | chat | 17.7 | 28.33 | 48.0 | 6.0 | 32.7 | 11.0 | |
| tinyagent-7b | chat | 23.1 | 41.3 | 28.0 | 8.0 | 58.7 | 12.0 |
4 Experiments
Our evaluation framework rigorously tests intelligent agents in six key domains to ensure their readiness for diverse real-world challenges Ross et al. (2023). These areas include seamless LLM integration into OS with an emphasis on security and user interaction; proficiency in real DB operations using SQL Halevy et al. (2004); task execution on the simulated e-commerce platform WebShop(WS) Yao et al. (2022); constructing and using KGs for enhanced semantic understanding; employing the M2W dataset for complex web tasks, marking the first dataset for developing general web agents following language instructions; and applying abstract reasoning and visual tasks in the text-based ALFWorld(ALF) Shridhar et al. (2021). For more implementation and evaluation details, see Appendices A and B.
4.1 Dataset
The dataset for our research was meticulously constructed to comprehensively evaluate the capabilities of agents Gou et al. (2020). It was established through self-collected methods, aimed at providing a rich and diverse testing environment to thoroughly assess the performance of deep learning models across various tasks Sachdeva and McAuley (2023). The construction of the dataset included key processes such as data collection, filtering, enhancement, and knowledge distillation Chen and Liu (2018). Through detailed screening and processing, we ensured the accuracy and consistency of the dataset, retaining only high-quality samples directly related to the testing objectives Sachdeva and McAuley (2023). Faced with issues of data imbalance and insufficient samples, we utilized data augmentation and knowledge distillation techniques. Knowledge distillation helped us to extract the most valuable and representative information from the vast amount of collected data, thus building an efficient and refined testing dataset. This process significantly improved the quality and applicability of the dataset, providing a solid foundation for evaluating the capabilities of model agents Mishra and Marr (2017).
4.2 Evaluating Code Correction
As shown in the Table 1, in this study, we conducted a comprehensive performance evaluation of TinyAgent-1.8B and the CodeLlama series models (CodeLlama7B and CodeLlama13B), aiming to explore their multi-task checking capabilities, including but not limited to code correction, OS configuration, DB query optimization, and WS. The experimental results showed that TinyAgent-1.8B demonstrated a significant advantage in cross-task performance evaluation compared to the CodeLlama series models. This performance was not only significant in code correction tasks but also prominent in other checking tasks such as OS configuration, DB query optimization, and WS management. These findings highlight that TinyAgent-1.8B not only possesses efficient code analysis capabilities but is also widely applicable to the inspection and optimization of other complex systems.
4.3 Baselines
In the baseline section of our study, we’ve selected Qwen-1.8B and CodeLlama-7B as pivotal benchmarks to assess the TinyAgent series’ performance, excluding the CMAT framework’s influence.
| Models | OS | DB | KG | ALF | WS | M2W |
| tinyagent-7b | 27.3 | 43.0 | 38.0 | 10.0 | 61.8 | 14.0 |
| - agent only | 20.1 | 39.3 | 25.0 | 2.0 | 55.7 | 7.0 |
| - general only | 9.7 | 5.4 | 0.0 | 0.0 | 26.6 | 5.0 |
4.4 Results analysis
The results in Table 2 underscore the effectiveness of our fine-tuning methods, especially for the TinyAgent models. Tinyagent-1.8B demonstrates significant performance in the KG task, on par with advanced models like GPT-3.5. Tinyagent-7B also showcases its strengths, notably in the DB task, where it surpasses its foundational model Antonello et al. (2020), CodeLlama-7B, and offers competitive scores against GPT-4. These findings indicate the TinyAgent models’ capacity to match or even surpass models with larger parameters in certain aspects. Moreover, the CMAT framework’s potential to enhance the capabilities of smaller-scale models is highlighted, allowing the TinyAgent models to closely compete with the performance of advanced models such as GPT-4.
As illustrated in Figure 1, Our comparative analysis indicates that Tinyagent models, refined from Qwen-1.8B and CodeLlama-7B, exhibit superior performance to their base models. The incorporation of the CMAT framework further amplifies their functionality, equipping these small Models to match the capabilities of GPT-3.5. This performance boost is credited to CMAT’s optimization of model interactions and its strategic use of memory modes for specific tasks, confirming its effectiveness in enhancing the sophistication of fine-tuned models Deshpande et al. (2021).
4.5 Error analysis
In our testing framework’s error analysis, we observed common challenges in DB tasks faced by models, such as difficulties in understanding user requests, executing actions, and pre-action problem analysis. Many models simply respond with "OK" to specific instructions without performing actual SQL operations, indicating a gap in transforming user requests into database actions. Models often provide superficial acknowledgments without delivering precise execution or in-depth problem analysis, failing to meet user expectations. In contrast, the TinyAgent series excels in understanding and converting user requests into actual SQL operations, effectively comprehending and executing tasks. It provides clear responses and adheres to user-specified SQL formats, fulfilling user expectations comprehensively. Additionally, TinyAgent’s thorough pre-action problem analysis and reflection demonstrate its advanced problem-solving skills and deep understanding of issues.
4.6 Ablation Study
The Table 3 presents an ablation study on the TinyAgent-7B model, delineating the impact of agent-specific and general instructions on task performance. The composite model, TinyAgent-7B, demonstrates the highest efficacy, notably in WS and DB tasks, which implies its adeptness in handling complex e-commerce interactions and database management. The agent-only variant exhibits a decline in performance, suggesting that while task-specific instructions are crucial, they are not wholly sufficient for the breadth of tasks such as KG. The general-only model’s performance is considerably reduced across all tasks, with a complete inability to perform in KG and ALF, highlighting the indispensability of agent-specific instructions. This data underscores the necessity of integrating both agent-specific and general instructions to enhance the versatility and effectiveness of AI models in diverse task domains.
5 Conclusions
The main findings of our work reveal that carefully trained small-parameter models on excellent datasets can achieve performance comparable to that of large-parameter models. With the application of the CMAT framework, we further demonstrate the significant potential for performance improvement in large-parameter models, highlighting the importance of model design and optimization strategies for parameter size. In our evaluation, although most open-source LLMs performed poorly compared to API-provided models without optimization, some models displayed similar capabilities to API models after meticulous fine-tuning of the TinyAgent model. This finding emphasizes not only the importance of parameter size in handling real-world environmental interactions but also showcases the enormous potential of even smaller models through the CMAT framework and precise adjustment strategies.
Limitations
In this study, we demonstrated the potential for performance improvement by applying the CMAT framework to TinyAgent series models and other large language models (LLMs). However, there are clear limitations to the research: First, although most models showed improved performance, some models saw limited improvement due to weaker base agent capabilities, indicating that the effectiveness of the CMAT framework might vary significantly between different models; second, the limitations of datasets and task types could affect the broad applicability of the conclusions, while low-quality datasets could negatively impact model performance; lastly, although evaluations based on AgentBench ensured fairness, they might not fully reflect the complexity of real-world scenarios, and due to computational resource constraints, larger-scale models could not be tested. This underscores the importance of future work to consider a wider range of models, datasets, and task types, especially the implementation of optimization strategies and framework applications in resource-constrained situations.
Appendix A Implementation Settings
In this paper, we describe an experiment conducted using the Low-Rank Adaptation (LoRA) fine-tuning method to enhance the performance of various models Wang et al. (2005). The accuracy of the LoRA method is of paramount importance in dealing with personalized and emotionally rich content. It enables the models to adapt to new data features while maintaining their core capabilities Bai et al. (2015).
During the experiment, we set the temperature parameter of the test models to 0.7 to increase the diversity of the content, and adjusted the top-p value to 0.95 to improve the precision of the generated content. We employed a learning rate of 2e-4 and beta values of (0.9, 0.999) to ensure the stability of the training process. The batch size was set to 4, with gradient accumulation, to ensure efficiency within the limits of computational resources. To balance innovation and coherence, we used LoRA parameters with a rank of 8 and an alpha value of 32, and adjusted both the top-p value and the temperature parameter to 0.7.These adjustments significantly enhanced the models’ flexibility and accuracy in handling personalized and emotionally rich content.
| OS | DB | KG | ALF | WS | W2M | |
| Completed | 84.7 | 84.0 | 25.0 | 2.0 | 93.5 | 57.0 |
| CLE | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| Invalid Format | 0.0 | 3.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| Invalid Action | 0.0 | 0.0 | 0.0 | 96.0 | 0.0 | 8.0 |
| TLE | 15.3 | 13.0 | 75.0 | 2.0 | 6.5 | 35.0 |
Appendix B Evaluation criteria
Operating systems Integrating LLMs into operating systems offers vast potential for automating and optimizing tasks. This integration demands a secure, user-friendly interface for effective LLM-OS interaction and requires LLMs to accurately understand the OS context for informed operations. Ensuring the safety of these operations is paramount to prevent misuse. Moreover, the system must handle errors and provide clear feedback to users, enhancing interaction and control. Addressing these aspects can revolutionize computer interaction and efficiency across industries.
Database Database (DB). Due to the crucial and challenging nature of database analysis in many daily affairs, it is paramount to examine the abilities of LLMs to operate on real databases via SQL. Previous research has placed significant emphasis on individual procedures, such as showcasing the effectiveness of LLMs in automating database access through T5QL, a new SQL generation method. Additionally, utilizing fine-tuned LLMs (such as GPT-3.5) to extract and link complex scientific information from scientific texts has demonstrated the capacity of LLMs to obtain structured knowledge from unstructured text and subsequently construct large databases Dunn et al. (2022).
WebShop represents an innovative simulation of an e-commerce website environment, featuring 1.18 million real-world products and 12,087 crowd-sourced text instructions. This platform challenges agents to navigate through multiple types of webpages and perform a variety of actions to find, customize, and purchase products according to given instructions. WebShop presents several challenges, including understanding compositional instructions, query (re-)formulation, dealing with noisy text in webpages, and conducting strategic exploration.
Knowledge Graphs The utilization of LLMs in constructing and interacting with knowledge graphs (KG) offers a promising avenue for enhancing semantic understanding and information retrieval. This involves assessing the models’ ability to not only generate but also interpret complex interrelations within data, facilitating more intuitive and context-aware responses. The effectiveness of LLMs in this domain could significantly improve AI’s capacity for reasoning and decision-making based on structured knowledge.
Mind2Web Mind2Web (M2W) is a dataset for developing web agents that perform complex tasks on real websites via language instructions. It features over 2,000 tasks across 137 sites from 31 domains. M2W’s real web environments and diverse user interactions make it a crucial platform for advancing AI navigation capabilities.
ALFWorld bridges interactive TextWorld environments with embodied tasks from the ALFRED dataset, enabling agents to learn abstract strategies and apply them to real-world tasks. It facilitates abstract reasoning and concrete execution, allowing agents to plan actions in a text-based simulator and then execute these tasks in a visual environment. This approach enhances agent generalization and problem-solving skills across various domains, such as language understanding and visual navigation, by leveraging a modular design that simplifies research improvements.
Appendix C Details about Experiment
| Evaluation Method | BLEU-4 | ROUGE-1 | ROUGE-2 | ROUGE-L |
| prompt - High-quality | 44.4 | 57.3 | 35.0 | 42.5 |
| prompt - Low-quality | 15.2 | 27.4 | 10.3 | 16.8 |
| without prompts | 26.8 | 47.2 | 30.2 | 26.7 |
In our research, we focused on exploring the impact of meticulously crafted prompts on the performance of Natural Language Processing (NLP) models in code correction tasks as shown in Table 5. By examining prompts of varying quality—high-quality, low-quality, and no prompts—we aimed to reveal how these factors influence the model’s capability, especially in identifying and correcting programming errors. Our findings indicate that fine-tuned high-quality prompts significantly enhance the model’s ability to write and rigorously check code, enabling it to produce high-quality code and conduct stricter code inspections. As demonstrated in the Table 23 25 .In contrast, low-quality prompts might lead to the model generating nonsensical and ineffective code, and in situations where correct code is provided, the model might even alter it to incorrect code. The decline in model output quality in these instances is primarily due to the vague and unclear descriptions provided by the prompts and the lack of specific guidance, which prevents the model from effectively utilizing the knowledge it acquired during training to focus on key error points and make accurate correction decisions. Moreover, our study also showed that in the absence of any prompt guidance, the model’s performance tends to be limited by the quality and scope of its training data, particularly when faced with novel or complex error types. This finding underscores the importance of high-quality prompts in designing effective NLP models for complex tasks like code correction. High-quality prompts not only guide the model to more accurately understand and address specific programming issues but also reduce noise in the processing, thereby improving overall efficiency and accuracy.
Appendix D Prompts for Tasks
Appendix E Examples of Tasks
All tasks in the dataset are divided into six categories. For each category, we present an example of successful execution as shown in Table 13 22.
DATABASE(DB) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
KNOWLEDGE GRAPH(KG) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
KNOWLEDGE GRAPH(KG) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
WEBSHOP(WS) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
ALFWORLD(ALF) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
MIND2WEB(M2W) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
OPERATING SYSTEM(OS) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
OPERATING SYSTEM(OS) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
KNOWLEDGE GRAPH(KG) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
KNOWLEDGE GRAPH(KG) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
ALFWORLD(ALF) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
MIND2WEB(M2W) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
MIND2WEB(M2W) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
WEBSHOP(WS) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
WEBSHOP(WS) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
DATABASE(DB) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
DATABASE(DB) {mdframed}[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]