### v6

# CMAT: A Multi-Agent Collaboration Tuning Framework for Enhancing Small Language Models

Abstract
Open large language models (LLMs) have significantly advanced the field of natural language processing, showcasing impressive performance across various tasks. Despite the significant advancements in LLMs, their effective operation still relies heavily on human input to accurately guide the dialogue flow, with agent tuning being a crucial optimization technique that involves human adjustments to the model for better response to such guidance. Addressing this dependency, our work introduces the TinyAgent model, trained on a meticulously curated high-quality dataset. We also present the Collaborative Multi-Agent Tuning (CMAT) framework, an innovative system designed to augment language agent capabilities through adaptive weight updates based on environmental feedback. This framework fosters collaborative learning and real-time adaptation among multiple intelligent agents, enhancing their context-awareness and long-term memory. In this research, we propose a new communication agent framework that integrates multi-agent systems with environmental feedback mechanisms, offering a scalable method to explore cooperative behaviors. Notably, our TinyAgent-7B model exhibits performance on par with GPT-3.5, despite having fewer parameters, signifying a substantial improvement in the efficiency and effectiveness of LLMs.
Index Terms:
Collaborative Multi-Agent Framework, Language Model Fine-Tuning, Reinforcement Learning from Human FeedbackI Introduction
In the rapid development of the field of artificial intelligence, large language models (LLMs) such as BERT and GPT-4 [1] have become important cornerstones of natural language processing (NLP). These models utilize the Transformer architecture and effectively capture long-distance dependencies through multi-head self-attention mechanisms, demonstrating strong capabilities across various NLP tasks. With technological advancements, the performance and application scope of LLMs continue to expand, promising significant improvements in computational efficiency and functionality, including anticipated advanced features such as self-improvement, self-checking, and sparse expert models [2].
However, it is noteworthy that the success of these models largely depends on human input to guide the correct dialogue. This dependency requires users to provide relevant and precise prompts based on their intentions and the feedback from the chat agent, raising a critical question: Can we replace human intervention with autonomous communication agents capable of steering conversations towards task completion with minimal human supervision?
Our research is driven by the need to overcome the significant challenges faced by LLMs in real-world deployments, particularly the high computational resource requirements, data biases, and lack of robustness. These issues limit their applicability in resource-constrained environments and highlight the urgency of enhancing model efficiency and adaptability [5, 6]. As demonstrated by Figure 1, we aim to address these limitations by optimizing models and training methods to enable smaller models to match the performance levels of larger models. Additionally, recognizing the potential of MAS to improve processing efficiency and system adaptability through agent cooperation, we seek to develop a collaborative agent framework. This framework aims to facilitate effective cooperation among agents, thereby overcoming the performance gap and propelling further research and innovation in the field of LLMs [7, 8]. In our experiments, we evaluated the capabilities of large models with and without the use of prompts and observed that low-quality prompts can significantly degrade model performance. Consequently, we propose the Collaborative Multi-Agent Tuning (CMAT) framework.
The CMAT framework introduces a structured environment where individual agents, each with specialized roles and capabilities, work together to process information, make decisions, and solve complex tasks [9]. By sharing insights and learning from interactions within this multi-agent ecosystem, the framework allows for a more scalable and flexible approach to training LLMs [10]. This collaborative effort not only helps in bridging the gap in performance between smaller and larger models but also fosters a more resilient system capable of adapting to new challenges without extensive human intervention [11]. Through CMAT, we aim to push the boundaries of what is possible with LLMs, making them more accessible and effective for a wider range of applications [12].
The main contributions of our work are as follows:
We propose the CMAT framework which represents an innovative approach that allows for dynamic and real-time memory updates within multi-agent systems.
We design a novel role-playing mechanism for precise task allocation and enhanced agent communication, significantly boosting overall performance and cooperation.
We evaluated the fine-tuned TinyAgent models across multiple agent tasks, finding that in certain scenarios, their performance rivals that of advanced LLMs like GPT-4 and agentlm [13], demonstrating the potential efficiency and capabilities of compact models.
II Related Work
II-A LLMs Applications in a Multi-Agent Framework
We explore the applications of LLMs within multi-agent systems, highlighting their role versatility as users, assistants, and checkers, and their capability to offer bespoke support and solutions across such environments [14, 15]. LLMs showcase remarkable adaptability to tasks through methods like supervised fine-tuning and real-time feedback learning, notably in tasks that require a sophisticated understanding and execution related to operating systems or databases [16, 17]. Furthermore, LLMs are adept at enhancing communication and collaboration among agents, a critical component for addressing complex issues that necessitate multi-role coordination [18]. Nevertheless, LLMs encounter specific challenges within multi-agent frameworks, especially in situations that demand a nuanced contextual comprehension and sustained memory retention, as well as adapting to fast-evolving environments and unforeseeable tasks [19]. Issues such as data bias, security concerns, and the intricacies of crafting effective protocols for multi-agent cooperation stand as significant hurdles in this domain [20, 21]. Thus, by summarizing LLMs’ roles in multi-agent frameworks, we underscore the critical need for continued innovation and research exploration, aimed at overcoming these technological hurdles and leveraging the full potential of LLMs in complex systems [22].
To enhance the adaptability and collaborative capabilities of LLMs in multi-agent systems, we’ve implemented memory modes, including long-term support and short-term memory with environmental feedback [23]. This allows LLMs to better interact, learn, and adapt in dynamic environments, leveraging past experiences and responding to changes swiftly.
II-B The tuning method for LLMs
The main tuning methods include supervised fine-tuning and reinforcement learning [24]. Supervised fine-tuning enhances performance by training models on specific task datasets, and is especially suitable for tasks such as natural language understanding (NLU) [25]. On the other hand, reinforcement learning, guided by reward mechanisms, is suitable for handling complex and variable tasks [26]. The effective combination of these two methods can significantly improve the performance of LLMs in various tasks. Notably, LLMs of reduced scale, such as those encompassing 1.8 billion parameters, can achieve performance levels akin to those of models with greater parameter counts, like 6 billion parameters, when supported by high-quality datasets [27]. This demonstrates that excellent data quality and appropriate tuning strategies play a decisive role in the performance of LLMs. Therefore, investing efforts in improving data quality and choosing the right tuning methods is essential for achieving optimal performance of LLMs in various application scenarios [25]. Through our work combining supervised fine-tuning with reinforcement learning, we’ve notably advanced LLM performance across a spectrum of tasks, showcasing significant improvements in task-specific benchmarks [24].
III Methodology
The Collaborative Multi-Agent Language Model Tuning (CMAT) framework improves decision-making, controllability, and efficiency in complex systems by coordinating three roles: User (), Assistant (), and Checker (). We fine-tune language models using LoRA [28], P-Tuning [29], and integrate ideas from RLHF [30], Chain of Thought (CoT), and ReAct to enhance reasoning and adaptability.
III-A Agent Roles and Actor-Critic Dynamics
In CMAT, the User () provides inputs, the Assistant () acts as the Actor generating actions, and the Checker () serves as the Critic providing feedback. At time , given input , the Assistant produces action based on its policy . The Checker evaluates this action and returns feedback , guiding continuous policy refinement.
III-B Learning Strategy
III-B1 Supervised Fine-Tuning
We start by fine-tuning the Assistant model on a labeled dataset :
where is the cross-entropy loss. This step ensures the model can produce useful actions before incorporating feedback loops.
III-B2 Incorporating CoT and ReAct
To enhance reasoning, we adopt Chain of Thought (CoT) to generate intermediate reasoning steps before the final action:
Additionally, ReAct interleaves reasoning and acting tokens, encouraging more accurate and context-aware decisions.
Empirically, generating reasoning steps (CoT) first, then the final answer, improves correctness. This suggests that explicitly formulating the thought process guides the model towards better actions.
III-B3 Feedback-Driven Policy Optimization
After supervised fine-tuning, the Assistant refines its policy using Checker feedback. We adopt an Actor-Critic-like scheme without full RL exploration. The Assistant updates its parameters via:
where is the error term:
Here, is derived from feedback . The Checker updates its value function:
Unlike traditional RL that involves environment exploration, here the process relies on direct Checker feedback. The iterative cycle refines the Assistant’s decision-making, ensuring alignment with the desired standards.
III-C Checker-In-The-Loop Mechanism
By positioning the Checker directly in the training loop, we ensure that the Assistant receives immediate corrective signals whenever it deviates from expected behavior
III-D Memory Management and Self-Reflection
The Assistant maintains two memories: short-term and long-term . Short-term memory tracks recent context:
while long-term memory stores significant insights via self-reflection:
Using the reflection, the Assistant updates its parameters:
where encodes feedback-driven loss and incorporates improvements identified via reflection.
IV Experiments
Our evaluation framework rigorously tests intelligent agents in six key domains to ensure their readiness for diverse real-world challenges [31]. These areas include seamless LLM integration into OS with an emphasis on security and user interaction; proficiency in real DB operations using SQL [32]; task execution on the simulated e-commerce platform WebShop(WS) [33]; constructing and using KGs for enhanced semantic understanding; employing the M2W dataset for complex web tasks, marking the first dataset for developing general web agents following language instructions; and applying abstract reasoning and visual tasks in the text-based ALFWorld(ALF) [34].
IV-A Dataset
The dataset for our research was meticulously constructed to comprehensively evaluate the capabilities of agents [35]. It was established through self-collected methods, aimed at providing a rich and diverse testing environment to thoroughly assess the performance of deep learning models across various tasks [36]. The construction of the dataset included key processes such as data collection, filtering, enhancement, and knowledge distillation [37]. Through detailed screening and processing, we ensured the accuracy and consistency of the dataset, retaining only high-quality samples directly related to the testing objectives [36]. Faced with issues of data imbalance and insufficient samples, we utilized data augmentation and knowledge distillation techniques. Knowledge distillation helped us to extract the most valuable and representative information from the vast amount of collected data, thus building an efficient and refined testing dataset. This process significantly improved the quality and applicability of the dataset, providing a solid foundation for evaluating the capabilities of model agents [38].
IV-B Evaluating Code Correction
As shown in the Table I, in this study, we conducted a comprehensive performance evaluation of TinyAgent-1.8B and the CodeLlama series models (CodeLlama7B and CodeLlama13B), aiming to explore their multi-task checking capabilities, including but not limited to code correction, OS configuration, DB query optimization, and WS.
| Model | BLEU-4 | ROUGE-1 | ROUGE-2 | ROUGE-L |
| codellama-7b | 25.01 | 45.91 | 29.83 | 26.24 |
| codellama-13b | 26.96 | 45.31 | 29.54 | 25.91 |
| tinyllama-1.8b | 43.38 | 59.86 | 37.81 | 42.86 |
The experimental results showed that TinyAgent-1.8B demonstrated a significant advantage in cross-task performance evaluation compared to the CodeLlama series models. This performance was not only significant in code correction tasks but also prominent in other checking tasks such as OS configuration, DB query optimization, and WS management. These findings highlight that TinyAgent-1.8B not only possesses efficient code analysis capabilities but is also widely applicable to the inspection and optimization of other complex systems.
IV-C Baselines
In the baseline section of our study, we’ve selected Qwen-1.8B and CodeLlama-7B as pivotal benchmarks to assess the TinyAgent series’ performance, excluding the CMAT framework’s influence.
| LLM Type | Models | VER | OS | DB | KG | ALF | WS | M2W |
| API | gpt-3.5-turbo | 0613 | 31.6 | 15.7 | 25.9 | 16.0 | 64.1 | 16.0 |
| gpt-4 | 0613 | 42.4 | 32.0 | 58.8 | 78.0 | 61.6 | 29.0 | |
| text-davinci-003 | - | 20.1 | 16.3 | 34.9 | 20.0 | 61.7 | 26.0 | |
| text-davinci-002 | - | 8.3 | 16.7 | 41.5 | 16.0 | 56.3 | 9.0 | |
| OSS | tinyllama-1.1b [39] | - | 2.8 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| opt-1.3b [40] | - | 0.7 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | |
| opt-2.7b | - | 1.4 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | |
| qwen-1.8b | chat | 10.4 | 22.67 | 6.8 | 0.0 | 26.6 | 5.0 | |
| chatglm2-6b 111https://github.com/thudm/chatglm2-6b | v1.1 | 4.2 | 1.3 | 0.0 | 0.0 | 0.0 | 0.0 | |
| codellama-7b | instruct | 9.7 | 2.7 | 0.0 | 0.0 | 14.3 | 5.0 | |
| llama2-7b [41] | chat | 0.0 | 4.2 | 8.0 | 0.0 | 11.6 | 7.0 | |
| zephyr-7b [42] | alpha | 12.5 | 9.7 | 5.0 | 8.0 | 45.0 | 11.0 | |
| baichuan2-6b [43] | chat | 2.8 | 9.7 | 0.0 | 0.0 | 6.1 | 11.0 | |
| mpt-7b 222https://github.com/mosaicml/llm-foundry/ | chat | 5.6 | 9.7 | 12.7 | 0.0 | 0.0 | 0.0 | |
| qwen-7b | chat | 12.5 | 13.0 | 7.0 | 34.3 | 0.0 | 0.0 | |
| agentlm-7b | chat | 14.6 | 33.0 | 9.0 | 16.4 | 18.4 | 10.0 | |
| agentlm-7b(SFT) | chat | 17.4 | 37.0 | 10.0 | 17.4 | 26.6 | 10.0 | |
| tinyagent-1.8b | chat | 17.7 | 28.33 | 48.0 | 6.0 | 32.7 | 11.0 | |
| tinyagent-7b | chat | 23.1 | 41.3 | 28.0 | 8.0 | 58.7 | 12.0 |
IV-D Results analysis
The results in Table II underscore the effectiveness of our fine-tuning methods, especially for the TinyAgent models. Tinyagent-1.8B demonstrates significant performance in the KG task, on par with advanced models like GPT-3.5. Tinyagent-7B also showcases its strengths, notably in the DB task, where it surpasses its foundational model [44], CodeLlama-7B, and offers competitive scores against GPT-4. These findings indicate the TinyAgent models’ capacity to match or even surpass models with larger parameters in certain aspects. Moreover, the CMAT framework’s potential to enhance the capabilities of smaller-scale models is highlighted, allowing the TinyAgent models to closely compete with the performance of advanced models such as GPT-4.
As illustrated in Figure 1, Our comparative analysis indicates that Tinyagent models, refined from Qwen-1.8B and CodeLlama-7B, exhibit superior performance to their base models. The incorporation of the CMAT framework further amplifies their functionality, equipping these small Models to match the capabilities of GPT-3.5. This performance boost is credited to CMAT’s optimization of model interactions and its strategic use of memory modes for specific tasks, confirming its effectiveness in enhancing the sophistication of fine-tuned models [45].
Table III presents the impact of different prompting strategies on performance metrics. High-quality prompts significantly outperform low-quality prompts and scenarios without prompts across all evaluation metrics, demonstrating the importance of prompt design in optimizing model performance.
| Evaluation Method | BLEU-4 | ROUGE-1 | ROUGE-2 | ROUGE-L |
| prompt - High-quality | 44.4 | 57.3 | 35.0 | 42.5 |
| prompt - Low-quality | 15.2 | 27.4 | 10.3 | 16.8 |
| without prompts | 26.8 | 47.2 | 30.2 | 26.7 |
IV-E Error analysis
In our testing framework’s error analysis, we observed common challenges in DB tasks faced by models, such as difficulties in understanding user requests, executing actions, and pre-action problem analysis. Many models simply respond with "OK" to specific instructions without performing actual SQL operations, indicating a gap in transforming user requests into database actions. Models often provide superficial acknowledgments without delivering precise execution or in-depth problem analysis, failing to meet user expectations. In contrast, the TinyAgent series excels in understanding and converting user requests into actual SQL operations, effectively comprehending and executing tasks. It provides clear responses and adheres to user-specified SQL formats, fulfilling user expectations comprehensively. Additionally, TinyAgent’s thorough pre-action problem analysis and reflection demonstrate its advanced problem-solving skills and deep understanding of issues.
As illustrated in Table IV, the distribution of various execution results across six tasks highlights the prevalence of specific error types, such as exceeding task limits (TLE) and invalid actions, which point to limitations in LLM agents’ reasoning and decision-making within constrained timeframes.
| Execution Results | OS | DB | KG | ALF | WS | M2W |
| Completed | 84.7 | 84.0 | 25.0 | 2.0 | 93.5 | 57.0 |
| CLE | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| Invalid Format | 0.0 | 3.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| Invalid Action | 0.0 | 0.0 | 0.0 | 96.0 | 0.0 | 8.0 |
| TLE | 15.3 | 13.0 | 75.0 | 2.0 | 6.5 | 35.0 |
IV-F Ablation Study
The Table V presents an ablation study on the TinyAgent-7B model, delineating the impact of agent-specific and general instructions on task performance. The composite model, TinyAgent-7B, demonstrates the highest efficacy, notably in WS and DB tasks, which implies its adeptness in handling complex e-commerce interactions and database management. The agent-only variant exhibits a decline in performance, suggesting that while task-specific instructions are crucial, they are not wholly sufficient for the breadth of tasks such as KG. The general-only model’s performance is considerably reduced across all tasks, with a complete inability to perform in KG and ALF, highlighting the indispensability of agent-specific instructions. This data underscores the necessity of integrating both agent-specific and general instructions to enhance the versatility and effectiveness of AI models in diverse task domains.
| Models | OS | DB | KG | ALF | WS | M2W |
| tinyagent-7b | 27.3 | 43.0 | 38.0 | 10.0 | 61.8 | 14.0 |
| - agent only | 20.1 | 39.3 | 25.0 | 2.0 | 55.7 | 7.0 |
| - general only | 9.7 | 5.4 | 0.0 | 0.0 | 26.6 | 5.0 |
V Conclusions
The main findings of our work reveal that carefully trained small-parameter models on excellent datasets can achieve performance comparable to that of large-parameter models. With the application of the CMAT framework, we further demonstrate the significant potential for performance improvement in large-parameter models, highlighting the importance of model design and optimization strategies for parameter size. In our evaluation, although most open-source LLMs performed poorly compared to API-provided models without optimization, some models displayed similar capabilities to API models after meticulous fine-tuning of the TinyAgent model. This finding emphasizes not only the importance of parameter size in handling real-world environmental interactions but also showcases the enormous potential of even smaller models through the CMAT framework and precise adjustment strategies.